1) How many molecules (not moles) of NH3 are produced from 3.66×10−4 g of H2?

Express your answer numerically as the number of molecules.

* At least 27 μg of tetrahydrocannabinol (THC), the active ingredient in marijuana, is required to produce intoxication. The molecular formula of THC is C21H30O2.

You may want to reference (Pages 93 - 98) Section 3.4 while completing this problem.

2) How many moles of THC does this 27 μg represent?

Express your answer using two significant figures.

3) How many molecules?

Express your answer using two significant figures.
Answer #1

1) N2 + 3H2 2NH3

3 moles H2 produce 2 moles of NH3

6g H2 produce 2 moles NH3

3.66×10-4 g H2 produce 2×3.66×10-4/6 moles NH3

= 1.22×10-4 moles of NH3

Number of molecules = 1.22×10-4×6.023×1023 molecules

= 7.35×1019 molecules

2) Moles of THC = 27×10-6/314.45 moles

= 8.6×10-8 moles

3) Number of molecules = 8.6×10-8×6.022×1023

= 5.2×1016 molecules